OpenStudy (anonymous):

Calc help please: sin(xy)=(x^2)cosy, find y prime

OpenStudy (loser66):

how far are you?

OpenStudy (anonymous):

I've always had trouble deriving trigonometric functions.. I don't know where to start

OpenStudy (anonymous):

ok, I can get the right side actually. hold on

OpenStudy (anonymous):

-(x^2)siny + 2xcosy

OpenStudy (anonymous):

but what do I do with the sin(xy)?

OpenStudy (loser66):

yours = sin(xy ) - x^2cos(y) =0 now take derivative, term by term, first term sin(xy) = cos (xy) * (xy)' = cos (xy) [ y + xy']

OpenStudy (loser66):

got me?

OpenStudy (anonymous):

no still trying to understand

OpenStudy (anonymous):

yeah it just clicked before you posted that but thank you

OpenStudy (loser66):

the second term?? do it!!

OpenStudy (loser66):

hey, not finish yet, do the second term, please. You have a long process to get the final answer.

OpenStudy (anonymous):

I know I'm working it on paper

OpenStudy (anonymous):

cos(xy)(y+xy1) + (X^2)siny - 2xcosy

OpenStudy (loser66):

nup

OpenStudy (anonymous):

no?

OpenStudy (loser66):